# WeTok: Powerful Discrete Tokenization for High-Fidelity Visual Reconstruction

### v2

@@ Abstract
Visual tokenizer is a critical component for vision generation. However, the existing tokenizers often face unsatisfactory trade-off between compression ratios and reconstruction fidelity. To fill this gap, we introduce a powerful and concise WeTok tokenizer, which surpasses the previous leading tokenizers via two core innovations. (1) Group-wise lookup-free Quantization (GQ). We partition the latent features into groups, and perform lookup-free quantization for each group. As a result, GQ can efficiently overcome memory and computation limitations of prior tokenizers, while achieving a reconstruction breakthrough with more scalable codebooks. (2) Generative Decoding (GD). Different from prior tokenizers, we introduce a generative decoder with a prior of extra noise variable. In this case, GD can probabilistically model the distribution of visual data conditioned on discrete tokens, allowing WeTok to reconstruct visual details, especially at high compression ratios. Extensive experiments on mainstream benchmarks show superior performance of our WeTok. On the ImageNet 50k validation set, WeTok achieves a record-low zero-shot rFID (WeTok: $\mathbf{0.12}$ vs. FLUX-VAE: $\mathbf{0.18}$ vs. SD-VAE 3.5: $\mathbf{0.19}$) with a $\mathbf{400\%}$ compression ratio. Furthermore, our highest compression model achieves a zero-shot rFID of $3.49$ with a compression ratio of 768, outperforming Cosmos (384) $4.57$ which has only 50% compression rate of ours. Code and models are available: https://github.com/zhuangshaobin/WeTok.
Figure 1: Zero-shot reconstruction comparison with state-of-the-art tokenizers. (a) Our WeTok establishes a new state-of-the-art trade-off between compression and reconstruction performance among the compared methods. (b) WeTok achieves a significant improvement in reconstruction quality over previous discrete tokenizers such as VQ-VAE and Open-MAGVIT2.
@@ 1 Introduction
In visual generation, the high computational cost of pixel-based data is a central challenge (Chen et al., 2020b; Rombach et al., 2022b). Visual tokenizers are a key solution that uses an encoder to compress an image into a compact latent representation and a decoder to reconstruct it (Kingma & Welling, 2013; Rezende et al., 2014), allowing generative models to operate efficiently in this latent space (Rombach et al., 2022a). These tokenizers are broadly divided into two categories: continuous (Kingma & Welling, 2013) and discrete (Van Den Oord et al., 2017). Continuous tokenizers map images to a continuous latent space, while discrete tokenizers employ a quantizer to produce a finite set of codes. This architectural difference introduces a critical trade-off. Discrete tokenizers can achieve a higher compression ratio, but this efficiency often comes at the cost of lower reconstruction fidelity compared to continuous methods. This leads to a natural question: Can we build a discrete tokenizer that can maintain high compression as well as achieve high-fidelity reconstruction?
To achieve this goal, two critical issues must be resolved. (1) Scalable Codebook. To minimize quantization error of discrete tokenizers, the existing methods attempt to enlarge the codebook (Yu et al., 2024a; Zhao et al., 2024b; Sun et al., 2024). In particular, the Lookup-Free Quantization (LFQ) (Yu et al., 2024a) quantizes the latent features directly, which largely increases the codebook size for better reconstruction. However, the substantial memory and computational overhead required to manage such a large codebook during training hinders further scalability. (2) Generative Modeling. Discrete tokenizers are inherently deterministic. Rather than modeling data distribution of images, decoder is trained to reconstruct the expected value of images (Esser et al., 2020), corresponding to the latent codes from encoder. Such a manner is limited to capture rich diversity and fine details in the original images, leading to unsatisfactory reconstruction, particularly at high compression ratios.
To fill the gap, we inrtdouce a powerful discrete tokenizer, namely WeTok, as shown in Fig. 1 (b). Specifically, we introduce two concise designs to solve the issues above. First, we develop a Group-Wise Lookup-Free Quantization (GQ), which partitions the latent feature of each image into groups, and performs lookup-free quantization for each group. As shown in Tab. 2, our GQ addresses the challenge in LFQ where entropy loss (Chang et al., 2022; Jansen et al., 2019) causes memory usage to grow with the codebook size, while yielding superior reconstruction performance. Second, we introduce a Generative Decoder (GD), which mimics the GAN-style generator by adding a prior of extra noise variable. As shown in Fig. 7, GD can effectively model data distribution of input images, allowing to reconstruct visual details at high compression ratios.
@@ 2 Related Work
2.1 Continuous Tokenizer
Generative modeling in the pixel space typically requires extensive compute resources (Chen et al., 2020a; Ho et al., 2020). Subsequent works (Rombach et al., 2022b; Podell et al., 2023; Peebles & Xie, 2023; Esser et al., 2024b; Batifol et al., 2025; Zhuang et al., 2025; 2024; Esser et al., 2024a; Zha et al., 2025; Chen et al., 2025) adopt VAE (Kingma & Welling, 2013), which projects visual content from pixels to latent features, achieving efficient and photo-realistic visual generation at high resolution. FLUX-VAE (Batifol et al., 2025) shows the state-of-the-art performance in both reconstruction quality and generalization ability across all continuous tokenizers. However, continuous tokenizer is criticized for its low compression rate, because latent features are usually stored and calculated in float32 or bfloat16. Therefore, discrete tokenizers that can store data in int or bool seem to be more promising in terms of compression capabilities.
2.2 Discrete Tokenizer
Traditional works like VQVAE (Van Den Oord et al., 2017) and VQGAN (Esser et al., 2021) employ vector-quantization (VQ) to transform visual input into discrete tokens. But they both suffer from low reconstruction quality caused by quantization error and instability of the codebook utilization. To overcome these drawbacks, subsequent works have explored several paths. One line of work introduces specific optimization strategies or modules to improve performance (Lee et al., 2022b; Shi et al., 2024; Zhu et al., 2024; Yu et al., 2024b). Another line of work focuses on mitigating the training instability when scaling up the codebook size by using grouped codebooks (Ma et al., 2025; Jia et al., 2025; Zhang et al., 2025). These methods split the input feature into groups along the channel dimension, where each group is then looked up using a sub-codebook. However, despite these improvements, VQ-based tokenizers still introduce additional inference and training costs due to the lookup operation (Yu et al., 2021b; Lee et al., 2022b; Fang et al., 2025). MAGVIT-v2 (Yu et al., 2024a) introduces Lookup-Free Quantization (LFQ) to address this extra cost and proposes the entropy loss (Chang et al., 2022; Jansen et al., 2019) to ensure the utilization of the codebook. However, the entropy loss causes unaffordable memory cost as it scales linearly with the codebook size, limiting the further expansion of the codebook. Binary Spherical Quantization (BSQ) (Zhao et al., 2024a) is proposed to mitigate this memory issue by assuming independence between the bits of the binary code, while this strong assumption can lead to performance degradation compared to LFQ. Concurrent to the above work, the GQ in our WeTok does not rely on explicit codebooks, and can eliminate the memory usage caused by entropy loss while having better performance than LFQ.
2.3 Autoregressive Visual Generation
The autoregressive (AR) modeling paradigm, which underpins modern Large Language Models (LLMs) (Vaswani et al., 2017), has been successfully adapted for visual generation (Chen et al., 2020a), where models learn to predict sequences of discrete tokens for images (Ramesh et al., 2021b; Ding et al., 2021; Liu et al., 2024) and videos (Hong et al., 2022; Kondratyuk et al., 2023). Recent AR models (Sun et al., 2024; Team, 2024; Wu et al., 2025a; Wang et al., 2024b; Liu et al., 2025) achieve remarkable image quality, highlighting the significant potential of this paradigm. Notably, the success of AR models is critically dependent on the visual tokenizer. Therefore, we adopt our WeTok to the LlamaGen (Sun et al., 2024) framework to enable high-fidelity and efficient autoregressive generation. This shows that our WeTok is not only capable of compression, but its compressed features are also suitable for generative models.
@@ 3 Method
In this section, we first establish the necessary preliminaries for discrete tokenization. We then introduce the Group-Wise Lookup-Free Quantization (GQ) to unify Lookup-Free Quantization (LFQ) (Yu et al., 2024a) and Binary Spherical Quantization (BSQ) (Zhao et al., 2024b). Finally, we present a Generative Decoder (GD) specifically engineered for high-compression scenarios to reconstruct high-fidelity outputs from the compact representations generated by our GQ.
@@ 3.1 Preliminaries
Vector Quantized Variational Autoencoder (Esser et al., 2020). Given an image $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$, VQVAE first compresses it into latent feature $\mathcal{U}=\mathcal{E}(\mathcal{I})$, $\mathcal{U}\in\mathbb{R}^{h\times w\times d}$, through the encoder $\mathcal{E}$. Then it is quantized into latent codes $\mathcal{Q}$ by searching the nearest neighbor in the codebook $\mathcal{C}\in\mathbb{R}^{K\times d}=[\mathbf{c}_{1},\mathbf{c}_{2},\dots,\mathbf{c}_{K}]^{\top}$,
@@
Binary Spherical Quantization (Zhao et al., 2024b). When increasing the codebook size, i.e., increasing $d$, the $H(\cdot)$ calculation in the $\mathcal{L}_{\text{Entropy}}$ of LFQ will lead to significant memory consumption. To alleviate this issue, BSQ seeks to handle each binary of LFQ seperately. BSQ first rewrites the token entropy loss into binary seperated form as $\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}\sum_{k=1}^{d}H(q_{B}(\mathbf{c}[k]|\mathcal{U}[i,j,k]))$. Next, BSQ rewrites the codebook entropy loss as $\sum_{k=1}^{d}H(q_{B}(\mathbf{c}[k]|\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}\mathcal{U}[i,j,k]))$ by assuming the approximation $\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}q(\mathbf{c}|\mathcal{U}[i,j]))\approx\prod_{k=1}^{d}\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}q_{B}(\mathbf{c}[k]|\mathcal{U}[i,j,k]))$. Both operations reduce the variable space of the $H(\cdot)$ calculation from $\{-1,1\}^{d}$ to the linear combination of $d$ $\{-1,1\}$, significantly decreasing memory consumption. However, as the binary approximation for the codebook entropy loss proposed by BSQ would introduce additional errors, the performance of BSQ is inferior to that of LFQ under the same codebook size.
@@ 3.2 Group-Wise Lookup-Free Quantization
As shown in Eq. 4, the computational cost of $H(\cdot)$ increases linearly with the codebook size. To simultaneously address it and the optimization error issue of BSQ, we propose grouping the latent features first and then performing lookup-free quantization. As shown in Fig. 2, we group the latent features from the channel dimension, reshape $\mathcal{U}$ into $\mathcal{U}_{\text{G}}\in\mathbb{R}^{h\times w\times g\times d^{\prime}}$, where $d=gd^{\prime}$ and $g$ is a newly added dimension representing the number of groups and $d^{\prime}$ is group channel. For $k$-th group, there is a non-learnable grouped codebook $\mathcal{C}_{\text{GQ},k}=\{-1,1\}^{d^{\prime}}$. The conditional probability after grouping can be formulated as
@@
We transform the original token entropy loss calculation from the $\{-1,1\}^{d}$ space into grouped token entropy loss which is a linear combination of entropy from $g$ $\{-1,1\}^{d^{\prime}}$ spaces. This change eliminates the token entropy loss as the memory bottleneck.
However, for the codebook entropy loss, the $H(\sum\cdot)$ operation prevents us from leveraging the additivity of entropy to decompose the $\{-1,1\}^{d}$ space into a linear combination of multiple subspaces. We propose the assumption that
@@
Thus, we leverage the additivity of entropy to transform the codebook entropy loss into
@@
Similar to the token entropy loss, we transform the calculation of the codebook entropy loss from the $\{-1,1\}^{d}$ space into a grouped codebook entropy loss, which becomes a linear combination of entropy from $g$ $\{-1,1\}^{d^{\prime}}$ spaces.
Governed by the group number $g$, GQ provides a tunable trade-off between approximation accuracy and memory cost. When $g=1$, GQ degenerates to the LFQ formulation, which uses no approximation and thus achieves high precision at the cost of substantial memory consumption. Conversely, when $g=d$, GQ mirrors the BSQ approach, which employs a strong approximation to minimize memory usage but consequently suffers from significant optimization error. Experiments in Sec. 4.2 demonstrate that by selecting an appropriate intermediate value for $g$, GQ can significantly reduce memory overhead while introducing only minimal optimization error. This allows GQ to surpass even the standard LFQ baseline. Furthermore, this tunable design provides the flexibility to effectively scale the codebook to a virtually unlimited size.
3.3 Generative Decoder
where $\mathcal{D}$ is the discriminator. However, the GAN loss only serves to assist in improving the perceptual quality. In scenarios with high compression rates, the correspondence between $\mathcal{U}_{\mathcal{Q}}$ and the ground truth is likely not unique. In such cases, we need the decoder to become a generative model. As shown in Fig. 2, we randomly sample $\mathbf{z}\in\mathcal{N}(\mathbf{0},\mathbf{I})$, concatenate it with $\mathcal{U}_{\mathcal{Q}}$ along the channel dimension, and then feed the result into the decoder. In this way, the GAN loss is subsequently reformulated as
where $\mathcal{U}_{\mathcal{Q}}$ serves as a condition for the generation process. The development from Eq. 9 to Eq. 10 is not merely concatenating $\mathbf{z}$ to the input of the decoder. More importantly, the learning objective of the decoder shifts to modeling the transformation from Gaussian noise conditioned on $\mathcal{U}_{\mathcal{Q}}$ to the ground truth distribution, turning it into a GD.
@@ 3.3 Generative Decoder
To ensure training stability, we employ a two-stage training process. In the first stage, we train our WeTok with the reconstruction loss, i.e., Eq. 2, 6 and 8. In the second stage, we adapt the model for generative tasks. Specifically, we expand the channel dimension of the conv_in layer in decoder to accept z as additional input. To preserve the powerful reconstruction capabilities learned in the first stage, the newly added weights for these channels are zero-initialized. This strategy ensures that at the beginning of the second stage, the decoder’s behavior is identical to its pre-trained state. Consequently, this versatile approach allows our GD to be effectively integrated with various pre-trained discrete tokenizers that share a similar architectural foundation.
4 Experiments
Figure 8: Parameter ablation experiments of the Autoregressive model based on WeTok. Compared with Open-MAGVIT2, the performance of the AR model based on WeTok improves more significantly with the increase in the number of parameters.
Table 1: Memory usage ablation. We set $d^{\prime}$==88 in GFQ. OOM refers to out of memory.

### v3

@@ Abstract
Visual tokenizer is a critical component for vision generation. However, the existing tokenizers often face unsatisfactory trade-off between compression ratios and reconstruction fidelity. To fill this gap, we introduce a powerful and concise WeTok tokenizer, which surpasses the previous leading tokenizers via two core innovations. (1) Group-wise lookup-free Quantization (GQ). We partition the latent features into groups, and perform lookup-free quantization for each group. As a result, GQ can efficiently overcome memory and computation limitations of prior tokenizers, while achieving a reconstruction breakthrough with more scalable codebooks. (2) Generative Decoder (GD). Different from prior tokenizers, we introduce a generative decoder with a prior of extra noise variable. In this case, GD can probabilistically model the distribution of visual data conditioned on discrete tokens, allowing WeTok to reconstruct visual details, especially at high compression ratio. On the ImageNet 50k validation set, at a high-fidelity setting, WeTok achieves a record-low zero-shot rFID of $\mathbf{0.12}$, outperforming leading continuous tokenizers like FLUX-VAE ($0.18$) and SD-VAE 3.5 ($0.19$) with $\mathbf{400\%}$ compression ratio. Furthermore, in a high-compression regime, WeTok achieves a zero-shot rFID of $3.49$ at a $\mathbf{768\times}$ compression ratio, substantially surpassing Cosmos, which scores $4.57$ at only $\mathbf{50\%}$ our compression ratio. Code and models are available: https://github.com/zhuangshaobin/WeTok.
Figure 1: Zero-shot reconstruction comparison with state-of-the-art tokenizers. (a) Our WeTok establishes a new state-of-the-art trade-off between compression and reconstruction performance among the compared methods. (b) WeTok achieves a significant improvement in reconstruction quality over previous discrete tokenizers such as VQVAE and Open-MAGVIT2.
@@ 1 Introduction
In visual generation, the high computational cost of pixel-based data is a central challenge (Chen et al., 2020b; Rombach et al., 2022b). Visual tokenizer is a key solution that uses an encoder to compress image into a compact latent representation and a decoder to reconstruct it (Kingma & Welling, 2013; Rezende et al., 2014), allowing generative models to operate efficiently in latent space (Rombach et al., 2022a). These tokenizers are broadly divided into two categories: continuous (Kingma & Welling, 2013) and discrete (Van Den Oord et al., 2017). Continuous tokenizers map images to a continuous latent space, while discrete tokenizers employ a quantizer to produce a finite set of codes. This architectural difference introduces a critical trade-off. Discrete tokenizers can achieve a higher compression ratio, but this efficiency often comes at the cost of lower reconstruction fidelity compared to continuous methods. This leads to a natural question: Can we build a discrete tokenizer that can maintain high compression as well as achieve high-fidelity reconstruction?
To achieve this goal, two critical issues must be resolved. (1) Scalable Codebook. To minimize quantization error of discrete tokenizers, the existing methods attempt to enlarge the codebook (Yu et al., 2024a; Zhao et al., 2024c; Sun et al., 2024). In particular, the Lookup-Free Quantization (LFQ) (Yu et al., 2024a) quantizes the latent features directly, which largely increases the codebook size for better reconstruction. However, the substantial memory and computational overhead required to manage such a large codebook during training hinders further scalability. (2) Generative Modeling. Discrete tokenizers are inherently deterministic. Rather than modeling data distribution of images, decoder is trained to reconstruct the expected value of images (Esser et al., 2020), corresponding to the latent codes from encoder. Such a manner is limited to capture rich diversity and fine details in the original images, leading to unsatisfactory reconstruction, particularly at high compression ratios.
To fill the gap, we introduce a powerful discrete tokenizer, WeTok, which consists of two concise designs to solve the issues above. First, we develop a Group-Wise Lookup-Free Quantization (GQ), which groups the latents and employs LFQ for each group. As shown in Tab. 2 and Fig. 3, GQ addresses the challenge in LFQ where entropy loss (Chang et al., 2022; Jansen et al., 2019) causes memory usage to grow with the codebook, while yielding superior reconstruction performance. Furthermore, we analyzed that the theoretical error caused by GQ is strictly smaller than that of BSQ. Second, we introduce the GAN-style generator into the decoder (GD). In Fig. 7, GD effectively models data distribution of images, allowing to reconstruct visual details at high compression ratios.
@@ 2 Related Work
VQVAE (Van Den Oord et al., 2017) and VQGAN (Esser et al., 2021) employ vector-quantization (VQ) to transform visual input into discrete tokens. But they both suffer from low reconstruction quality caused by instability of the codebook utilization. To overcome these drawbacks, one line of work introduces optimization strategies or modules to improve performance (Lee et al., 2022b; Shi et al., 2024; Zhu et al., 2024; Yu et al., 2024c). Another line of work focuses on scaling up the codebook size by grouping codebooks (Ma et al., 2025; Jia et al., 2025; Zhang et al., 2025; Bai et al., 2024). ImageFolder (Li et al., 2024), DualToken (Song et al., 2025) and TokenFlow (Qu et al., 2024) use multiple codebooks to assist in optimizing model understanding and generation capabilities. However, VQ-based tokenizers still introduce additional costs due to the lookup operation (Yu et al., 2021b; Lee et al., 2022b; Fang et al., 2025). MAGVIT-v2 (Yu et al., 2024a) introduces Lookup-Free Quantization (LFQ) to address extra cost and proposes the entropy loss (Chang et al., 2022; Jansen et al., 2019) to ensure the utilization of the codebook. BSQ (Zhao et al., 2024b) assumes independence between the bits of the binary code to eliminate unbearable computational overhead from entropy loss, while this assumption leads to performance degradation. In contrast, WeTok does not rely on explicit codebooks, and eliminate the memory usage caused by entropy loss while having better performance than LFQ. More related works could be found in Sup. B.
DiTo (Chen et al., 2025b), consistency decoder (Betker et al., ) and $\epsilon$-VAE (Zhao et al., 2024a) introduce diffusion into decoder to help visual reconstruction of continuous image tokenizer. Our WeTok is the first to introduce a generative decoder into discrete tokenizer, overcoming the instability of generative training caused by the stopping gradient estimation operation. Compare to diffusion modeling, Compared to diffusion modeling, our GAN-based WeTok achieves more efficient single-step sampling while also achieving state-of-the-art reconstruction performance.
@@ 3 Method
In this section, we first establish the necessary preliminaries for discrete tokenization. We then introduce the Group-Wise Lookup-Free Quantization (GQ) to unify Lookup-Free Quantization (LFQ) (Yu et al., 2024a) and Binary Spherical Quantization (BSQ) (Zhao et al., 2024c). Finally, we present a Generative Decoder (GD) specifically engineered for high-compression scenarios to reconstruct high-fidelity outputs from the compact representations generated by our GQ.
@@ 3.1 Preliminaries
[…]
@@
Binary Spherical Quantization (Zhao et al., 2024c). When increasing the codebook size, i.e., increasing $d$, the $H(\cdot)$ calculation in the $\mathcal{L}_{\text{Entropy}}$ of LFQ leads to significant memory consumption. To alleviate this issue, BSQ processes each binary of LFQ seperately. Firstly, rewrite the token entropy loss as $\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}\sum_{k=1}^{d}H(q_{B}(\mathbf{c}[k]|\mathcal{U}[i,j,k]))$. Next, the codebook entropy loss could be transformed into $\sum_{k=1}^{d}H(q_{B}(\mathbf{c}[k]|\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}\mathcal{U}[i,j,k]))$ by assuming the approximation $\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}q(\mathbf{c}|\mathcal{U}[i,j]))\approx\prod_{k=1}^{d}\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}q_{B}(\mathbf{c}[k]|\mathcal{U}[i,j,k]))$. Both operations reduce the variable space of the $H(\cdot)$ calculation from $\{-1,1\}^{d}$ to the linear combination of $d$ $\{-1,1\}$, significantly decreasing memory consumption. However, as the approximation for the codebook entropy loss would introduce additional errors, which leads to performance degradation.
@@ 3.2 Group-Wise Lookup-Free Quantization
[…]
@@
We first transform the token entropy loss from the $\{-1,1\}^{d}$ space into a linear combination of $g$ $\{-1,1\}^{d^{\prime}}$ spaces. This change eliminates the token entropy loss as the memory bottleneck.
For the codebook entropy loss, the $H(\sum\cdot)$ operation prevents us to decompose the $\{-1,1\}^{d}$ space into a linear combination of multiple subspaces. We propose the assumption that
@@
Thus, we could transform the codebook entropy loss into
@@
Similar to the token entropy loss, the codebook entropy loss is converted from $\{-1,1\}^{d}$ space into a grouped form, which eliminates the memory bottleneck caused by this term. GQ provides a tunable trade-off between approximation accuracy and memory cost by $g$. Experiments in Sec. 4.1 demonstrate that by selecting an appropriate $g$, GQ significantly reduces memory overhead while introducing minimal optimization error. This allows GQ to surpass the LFQ and BSQ. Furthermore, this tunable design provides the flexibility to scale the codebook to a virtually unlimited size.
Proposition 3.1.
For any choice of group $G$, the codebook entropy approximation error (as in Eq. 7) of our GQ method is smaller than that of the BSQ method.
Remark 3.2.
The detailed proof and definition of the approximation error of the Proposition 3.1 is deferred to Sup. A. We mainly adopt the order theory in abstract algebra to derive such a conclusion.
@@
3.3 Generative Decoder
where $\mathcal{D}$ is the discriminator. However, the GAN loss only serves to assist in improving the perceptual quality. In high compression ratio scenarios, the correspondence between $\mathcal{U}_{\mathcal{Q}}$ and the ground truth is likely not unique. In such cases, what we need is a generative decoder. As shown in Fig. 2, we randomly sample $\mathbf{z}\in\mathcal{N}(\mathbf{0},\mathbf{I})$, concatenate it with $\mathcal{U}_{\mathcal{Q}}$ along the channel dimension, and then feed the result into the decoder. In this way, the GAN loss is subsequently reformulated as
where $\mathcal{U}_{\mathcal{Q}}$ serves as a condition for the generation process. The development from Eq. 9 to Eq. 10 is not merely concatenating $\mathbf{z}$ to the input of the decoder. More importantly, the decoder shifts to modeling the transformation from Gaussian noise conditioned on $\mathcal{U}_{\mathcal{Q}}$ to the ground truth distribution. In previous discrete tokenizer, the decoder is trained to minimize a reconstruction loss like L2 or LPIPS. When a single, highly compressed discrete token $\mathcal{U}_{\mathcal{Q}}$ could correspond to multiple ground-truth images (e.g., different textures of fur , different patterns of leaves), the decoder learns to output the conditional expectation or the ”average” of all these possibilities. Our GD, by incorporating $\mathbf{z}$ as an additional input, transforms the decoder into a conditional generative model. It no longer learns a one-to-one mapping but instead models the conditional distribution. The noise vector $\mathbf{z}$ allows the decoder to sample one specific, plausible instance from this distribution. This single sample can contain coherent, high-frequency details (e.g., a specific fur texture), making the output appear much more realistic and less blurry than the ”average” image.
To ensure training stability, we employ a two-stage training strategy. In first stage, we train our WeTok with the reconstruction loss, i.e., Eq. 2, 6 and 8. In the second stage, we adapt the model for generative tasks. Specifically, we expand the channel dimension of the conv_in layer in decoder and employ zero-initialization to new channel to accept z as additional input. This strategy ensures that at the beginning of the second stage, the decoder’s behavior is identical to its pre-trained state.
Table 1: Memory usage ablation. We set $d^{\prime}$==88 in GQ. OOM refers to out of memory.